### v2

@@ Abstract
Outlier detection is an important data mining tool that becomes particularly challenging when dealing with nominal data. First and foremost, flagging observations as outlying requires a well-defined notion of nominal outlyingness. This paper presents a definition of nominal outlyingness and introduces a general framework for quantifying outlyingness of nominal data. The proposed framework makes use of ideas from the association rule mining literature and can be used for calculating scores that indicate how outlying a nominal observation is. Methods for determining the involved hyperparameter values are presented and the concepts of variable contributions and outlyingness depth are introduced, in an attempt to enhance interpretability of the results. The proposed framework is evaluated on both synthetic and real-world data sets, demonstrating comparable performance to state-of-the-art frequent pattern mining algorithms and even outperforming them in certain cases. The ideas presented can serve as a tool for assessing the degree to which an observation differs from the rest of the data, under the assumption of sequences of nominal levels having been generated from a Multinomial distribution with varying event probabilities.
@@ 1 Introduction
Outlier detection (also known as anomaly detection) is a crucial aspect of data analysis that has received increasing attention by data mining researchers. The wide range of uses of outlier detection in various domains, such as in cybersecurity [1, 18], sociology [39], healthcare [45] and finance [35], has driven the development of numerous algorithms designed to detect anomalies within a data set. Most of these algorithms are restricted to just one data type and that is continuous data. However, categorical data, that is binary, ordinal and nominal data [22], is prevalent in real-world data sets, yet very few outlier detection methods can handle this data type and authors do not always make the distinction between purely nominal and ordinal variables. Outliers in data sets with nominal variables may indicate a serious offence or a critical situation. For instance, unusual combinations of symptoms, diagnoses and demographic characteristics of patients could be indications of rare conditions in healthcare monitoring or uncommon combinations of claim types, policies and accident locations may hint fraud in insurance claims. This demands the development of outlier detection methods tailored for nominal data.
@@
One approach to outlier identification assuming an underlying discrete distribution, is extending ideas from the continuous to the discrete domain. This would typically require robust estimation of the parameters of some assumed underlying distribution, computation of robust distances and finally comparing the distances to a cutoff value [see 43, 37, 38, for examples of robust estimators]. The cutoff value is typically the $1-\alpha$ quantile of the distribution of the distances, assuming validity of the hypothesised data generating mechanism. In the discrete distribution setting, this is generally a hard task that this is not possible unless some model is assumed to have generated the counts in a contingency table. Log-linear models are widely used for this purpose and one can then infer what the expected proportions should be in a robust way [see for example 27, 28, 11]. However, these methods can become impractical when the number of nominal variables is very large and the analysis of a high-dimensional contingency table is required. The sparsity of the resulting contingency tables increases together with the dimensionality, which is something that needs to be accounted for as well.
@@
Another notable approach to outlier detection for nominal data is using proximity-based methods. These methods seek to detect outliers by generating a distance matrix between observations and finding the subjects that are furthest away from the rest of the points in the data set. Examples of such methods include [6] and [32]. However, distances in the categorical space often lack meaningful interpretation and implicitly make an equidistance assumption. As an example, the Hamming distance treats all mismatches equally. [42] introduced a distance metric for detecting outliers in categorical data; however, its application is restricted to $2\times 2$ contingency tables. Furthermore, these methods are very impractical in the presence of a large number of nominal variables, leading to a high computational cost. Information-theoretic methods have also been proposed for outlier identification of nominal data [30, 47, 48]. Such methods rely on some information-theoretic measure of data disorder, such as the entropy, and typically employ a local search algorithm to optimise an objective function. As a result, these methods rely heavily on heuristics and the optimisation process can be very costly. These problems are also common in recently proposed deep learning methods [for example 13, 14].
@@
One different approach to anomaly detection for nominal data is the rule-based approach that makes use of ideas from the association rule mining literature [4]. To the best of our knowledge, the first algorithm that adopted this mindset for dealing with the problem of outlier detection in nominal data was the Link-based Outlier and Anomaly Detection in Evolving Data sets (LOADED) algorithm of [19]. In fact, LOADED was developed for identifying anomalies in mixed-type data (that is for a combination of nominal and continuous features). An improvement of this algorithm in terms of computational cost was then presented by [36], with further ameliorations being proposed by [24] for the nominal outlier identification part. More precisely, [24] introduced the Outlier Detection for Mixed Attribute Datasets (ODMAD) algorithm. The primary concern with the way ODMAD deals with nominal features is that it treats outliers in the nominal domain in a way that disregards the dimension of the subspace in which these are found. Furthermore, the implementation of ODMAD involves certain threshold parameters, the values of which need to be defined by the user, without providing guidelines of how these could be chosen in a data-driven manner. In a similar fashion, several frequent pattern mining methodologies have been employed for quantifying nominal outlyingness, such as the Frequent Pattern Outlier Factor (FPOF) [21] and the Frequent Pattern Isolation (FPI) [26] algorithms. However, these methods also suffer from a lack of intuition regarding the choice of their associated hyperparameter values and do not take into account that longer sequences are inherently less frequent than their subsets. Moreover, they completely overlook whether an event is rare by nature. For instance, consider a rare disease that affects only a small proportion of the population but is known and not particularly noteworthy. Frequent pattern mining algorithms will most likely flag subjects suffering from that disease as potential outliers solely due to their low frequency within a data set of clinical records, even though these are completely normal and expected cases. This underscores the need for a framework that can incorporate such contextual information and produce results that align with the underlying ground truth.
@@
The rest of the paper is organised as follows: we define the concept of nominal outlyingness by adopting a rule-based mindset in Section 2 and we give a brief introduction to key concepts from the area of association rule mining. We then propose a score that quantifies abnormal behaviour in Section 3, focusing on the rationale behind its formulation and how the associated hyperparameters can be determined from the data by making the assumption of an underlying data generating mechanism. Some mathematical properties of our proposed score are included in Section 4. The efficacy of our method is illustrated via a series of simulations on synthetic and real-world data in Section 5, where our proposal is benchmarked against two state-of-the-art frequent pattern mining algorithms. We finally make some concluding remarks in Section 6.
@@ 2 Outlyingness in discrete space
In this Section we extend concepts of multivariate outlier detection for continuous data to discrete spaces. We define the concept of nominal outlyingness and give a brief introduction to basic notions from the association rule mining.
@@ 2.1 Definition of nominal outlyingness
One may find several definitions of outlyingness for continuous observations in the literature. All these definitions typically rely on some assumptions, for example related to the data generating mechanism [17, 7], or the location of each observation relative to the rest of the points in the feature space [e.g. 10, 33]. If nominal variables appear in a data set, using the location of points becomes difficult to work with and looking at distances in the categorical space is not ideal for reasons previously stated in Section 1. As we are dealing with point masses instead of densities, only a fixed number of values can occur, meaning that location may no longer be used as a criterion to assess anomalous behaviour. However, one can make use of the ideas related to the identification of outliers with respect to the mechanism that has been used to generate the data. In fact, all we have to assume is that the data has been generated from some underlying discrete distribution. Then, we turn our attention to observations of abnormally high or low mass with respect to this discrete distribution, just like we would suspect points of low density to be potential outliers in the continuous setting. This leads to two possible definitions of nominal outlyingness; one for observations whose nominal values are too frequent and one for too infrequent values, always with respect to an assumed discrete random variable. The notion of frequency is described in detail later in the section.
Notice that looking at both highly frequent and highly infrequent values is rather unsound; this is easy to realise by considering a toy example of one binary variable generated by a Bernoulli distribution for which the success probability has been significantly misspecified. In such a case, one of the two values will be treated as highly infrequent. However, this implicitly means that the other value will be highly frequent. Therefore, one needs to specify whether their interest lies in highly frequent or highly infrequent values, that is basically what they wish to define as a nominal outlier.
Extending these ideas to multiple nominal features requires taking combinations of their values into account. One potential downside of such an approach is that in the presence of multiple nominal variables taking a large number of values, it is very likely for almost every single combination of values to appear fewer times than expected. As a result, our definition of nominal outlyingness considers sequences of nominal values, weighing their contribution to outlyingness in a way that is inversely proportional to the sequence length. This is done in order to ensure that, if interest lies in highly infrequent sequences, combinations of values are seen with greater suspicion if they are shorter in length, as they are typically expected to be more abundant than their supersets in a data set. Based on these considerations, we can now define nominal outlyingness as follows:
@@ 2.2 Association rule mining background
Association rule mining is a method used to discover relationships of interest between sets of data items. The work of [4] was one of the first to focus on generating association rules between items in a large database of customer transactions and has paved the way for related concepts to be used in market basket analysis [46], customer segmentation [40] and other application areas. Below, we present some basic notions used in association rule mining. We later make use of these concepts in the development of a suitable score that quantifies how outlying a nominal observation is.
We start by defining an itemset as a set of items which occur together. Here, item refers to the value any variable takes for each observation. Since we are working with nominal data, an itemset is simply a sequence of categorical levels. We denote an itemset by $d$ and its length by $\lvert d\rvert$. Notice that $1\leq\lvert d\rvert\leq p$, where $p$ is the number of variables and $\lvert d\rvert=1$ refers to one level of a single nominal variable. Finally, we define the support of a an itemset $d$, denoted by $\operatorname{supp}(d)$, as the number of times the itemset $d$ appears in the data.
Given a large data set, some itemsets may appear more often than others. In an outlier identification problem, where the focus is on highly frequent or highly infrequent occurrences, determining whether an itemset is frequent or infrequent is of a major importance. An itemset can be considered as frequent when its support exceeds a minimum support threshold value, typically denoted by $\sigma$; otherwise it is infrequent. As we will see, this threshold value may differ according to the itemset that is being looked at, therefore we use the notation $\sigma_{d}$ instead, to denote that the threshold value is specific to an itemset $d$. Determining $\sigma_{d}$ is one of the most challenging problems in the association rule mining literature; this is something that we deal with later in the paper.
Finally, we introduce the Apriori principle and the related concept of support-based pruning. According to the Apriori principle [3], if an itemset is frequent, then all of its subsets must be frequent as well. Equivalently, if an itemset is infrequent, then all its supersets will also be infrequent. A direct implication of the Apriori principle is that one may not consider subsets or supersets of itemsets that are already found to be frequent or infrequent, respectively, something known as support-based pruning. The reason why support-based pruning is used is to avoid redundant computations and hence reduce the computational complexity of the search by trimming the space of itemsets that are being examined.
@@ 3 Proposed framework
The problem of outlier identification is generally complex in the sense that it is an unsupervised problem and thus, we can not be certain of whether an observation is truly outlying. However, what we can do is measure outlyingness with respect to a given scheme. By doing so, not only can we deduce how likely an observation is to be an outlier, but we can also compare observations in terms of their outlyingness. In this Section, we introduce a score that quantifies nominal outlyingness given Definition 1. We explain the rationale behind our score formulation and describe how values for the involved hyperparameters can be chosen. We finally introduce the concepts of nominal outlyingness depth and the contribution matrix which can be used for meta-analysis purposes.
@@ 3.1 Score of nominal outlyingness
Let us assume that we are working with a dataset $\mathcal{D}$ consisting of $p$ nominal variables $\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{p}$. We denote the number of levels for each variable $\boldsymbol{X}_{i}$ by $\ell_{i}$ and assume without loss of generality that its levels are encoded by the natural numbers $1,\ldots,\ell_{i}$ $(i=1,\ldots,p)$. Having defined some basic notions from the association rule mining literature, we treat each entry in a contingency table as an individual itemset. The total number of contingency tables is given by $12^{p}-$ 1 and the total number of itemsets that one can encounter is $\prod_{i=1}^{p}\left(\ell_{i}+1\right)-1$, assuming that an individual nominal feature $\boldsymbol{X}_{i}$ is a $(\ell_{i}\times 1)$-dimensional contingency table. We further define $\boldsymbol{\pi}_{i}$ to be the $\ell_{i}$-dimensional vector including the expected proportions for each level of $\boldsymbol{X}_{i}$ $(i=1,\ldots,p)$.
@@
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r}},$ (1)
@@
if we are interested in highly infrequent itemsets. Recall that $\operatorname{supp}(d)$ refers to the support of itemset $d$, as defined in Section 2. Otherwise, for highly frequent itemsets, the score of nominal outlyingness becomes:
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin[0,\sigma_{d}),\\ \lvert d\rvert\leq\operatorname*{\operatorname{MAXLEN}}\end{subarray}}\frac{\operatorname{supp}(d)}{\sigma_{d}\times\left(\operatorname*{\operatorname{MAXLEN}}-\lvert d\rvert+1\right)^{r}},$ (2)
@@
For the rest of this section, we will be assuming that nominal outliers are highly infrequent observations with respect to an underlying discrete distribution. Hence, our arguments will involve Expression (1); extending these to the case of highly frequent itemsets is straightforward.
The formulation in Expression (1) begins by looking at itemsets of unit length and compares their support to a lower threshold value $\sigma_{d}$. If the support lies outside $(\sigma_{d},n]$, the score of all points possessing that itemset is augmented in a way that is directly proportional to $\sigma_{d}$ and inversely proportional to the product of the itemset support and the itemset length raised to a non-negative power $r$. One can immediately notice the similarity of Expression (1) (with $r=1$ and a numerator of a unit) to the one used by [25] and [24]. The idea behind modifying the score to compute $\sigma_{d}/\operatorname{supp}(d)$ instead of just $1/\operatorname{supp}(d)$ is that we wish to ensure a more fair contribution of itemsets with respect to their expected probabilities. For instance, let us suppose that two highly infrequent itemsets $d_{1},d_{2}$ of unit length involving two distinct nominal variables, occur the same number of $\operatorname{supp}(d)$ times in a data set of $n$ observations, respectively. The nominal score of [24] would be the same for observations possessing $d_{1}$ and $d_{2}$. However, if the associated probabilities are given by $\pi_{d_{1}}<\pi_{d_{2}}$, one would expect the relevant threshold values to be $\sigma_{d_{1}}<\sigma_{d_{2}}$. Hence, it would make more sense to treat $d_{2}$ as more outlying compared to $d_{1}$, since the former was expected to occur much more frequently than the latter. Computing the ratio $\sigma_{d}/\operatorname{supp}(d)$ is a way to account for this and ensure that the score is augmented proportionally to the ratio of the minimum expected over the actual support of each highly infrequent itemset.
Our main contributions include a method for determining the support threshold values $\sigma_{d}$ that are specific to each itemset $d$. Moreover, we propose a suitable stopping criterion, that is a $\operatorname*{\operatorname{MAXLEN}}$ value, that accounts for the sparsity introduced in higher-dimensional contingency tables. We also derive some mathematical properties of $s(\cdot)$, such as when $s(\cdot)$ is maximised and what its maximum value is; these are included in Propositions 1 & 2 in Section 4.
3.2 Choice of hyperparameter values
The hyperparameters that are involved in the calculation of the score $s(\cdot)$ are the minimum support threshold $\sigma_{d}$, the maximum length of sequences considered $\operatorname*{\operatorname{MAXLEN}}$ and the exponent term $r$. We start by looking at the minimum support threshold value $\sigma_{d}$. This is defined with respect to the assumed underlying Multinomial distribution for each itemset. Assume without loss of generality that we are working with itemsets $d$ consisting of the first $\lvert d\rvert$ nominal variables. For such itemsets, the underlying Multinomial distribution that we consider is:
$\boldsymbol{X}=\boldsymbol{x}\mid d\sim\text{Multinomial}\left(1,\boldsymbol{p}^{d}\vcentcolon=\bigotimes\limits_{i=1}^{\lvert d\rvert}\boldsymbol{\pi}_{i}\right),$
where $\otimes$ is the Kronecker product. Notice how this formulation implicitly assumes independence of the nominal variables; a dependence structure between the nominal variables would not really change anything besides the definition of the probability vector, which would have to be explicitly provided by the user instead.
Returning to our original problem of determining a suitable threshold value for the support of an itemset $d$, we compute the support of the itemset and see if it falls within a specific range of reasonable values under the assumed model. This is done by constructing an upper one-sided $100(1-\alpha)\%$ confidence interval for the probability of observing itemset $d$. Since all itemsets of length $\lvert d\rvert$ form a Multinomial distribution with $\prod_{i=1}^{\lvert d\rvert}\ell_{i}$ possible outcomes, we use the method of [41] to construct $100(1-\alpha)\%$ simultaneous confidence intervals for the Multinomial proportions. The confidence intervals are constructed by approximating the Multinomial cumulative distribution function (CDF) using a theorem from [31]. The theorem approximates the Multinomial CDF as the product of Poisson probabilities multiplied by the probability mass function of a sum of truncated Poisson random variables, which is approximated using Edgeworth series. The two-sided $100(1-\alpha)\%$ confidence intervals for the components of the Multinomial proportions vector $\boldsymbol{p}^{d}$ are given by:
In a similar manner, the upper and lower one-sided $100(1-\alpha)\%$ confidence intervals are given by:
The confidence intervals obtained for the Multinomial proportions are used to determine what the least expected values for the itemset supports should be. Multiplying the lower bound of each upper one-sided interval by the sample size $n$ yields bounding values for the support of each itemset. These are in fact our values for $\sigma_{d}$, determining when an itemset is highly infrequent with respect to the underlying Multinomial distribution. This extends the idea of [36] and [24], who also augmented the nominal outlyingness scores of observations possessing highly infrequent itemsets using an expression similar to (1), but with a constant $\sigma$ value that was the same for all itemsets of any length.
As the itemset length grows, the associated contingency tables become more sparse. This leads to a serious issue for our procedure if one assumes that $p_{i}^{d}\overset{\lvert d\rvert\rightarrow p}{\longrightarrow}0$, since almost any itemset will be considered as highly infrequent. Besides, allowing for $\lvert d\rvert\rightarrow p$ leads to a substantial and potentially prohibitive increase in computational resources required. As a result, we seek to find a maximum allowed value for $\lvert d\rvert$, which we denote by $\operatorname*{\operatorname{MAXLEN}}$, so that itemsets of length up to $\operatorname*{\operatorname{MAXLEN}}$ will be included in our score computations. Determining $\operatorname*{\operatorname{MAXLEN}}$ so that it reflects on our concerns regarding sparsity can be done by requiring all upper threshold values $\sigma_{d}$ to be at least equal to two for all itemsets $d$. The rationale behind this is that if $\sigma_{d}<2\ \forall d:\lvert d\rvert=M$, then only itemsets of length $M\leq p$ with unit support will not be flagged as infrequent. This is a very restrictive case as most (if not all) observations will have their scores augmented due to being more frequent than expected. Thus, we define $\operatorname*{\operatorname{MAXLEN}}$ as follows:
[…]
The use of concepts from the association rule mining literature for the proposed score of nominal outlyingness presents a practical issue related to computational complexity. For a given value of $\operatorname*{\operatorname{MAXLEN}}$, the total number of contingency tables that can be generated is given by $\sum_{k=1}^{\operatorname*{\operatorname{MAXLEN}}}C^{p}_{k}$, where $C$ denotes the binomial coefficient. Moreover, the number of possible itemsets also gets very large which shows that under the presence of many variables, the total number of comparisons that need to be made can grow very rapidly, especially for large $\operatorname*{\operatorname{MAXLEN}}$. As a result, any computations that may be redundant should be omitted. Hence, we implement support-based pruning, excluding supersets of any highly infrequent itmsets from subsequent computations. That is, if $\lvert d_{1}\rvert<\operatorname*{\operatorname{MAXLEN}}$ and $\operatorname{supp}(d_{1})<\sigma_{d_{1}}$, then any itemset $d_{2}\supset d_{1}$ will not be checked in subsequent calculations.
Support-based pruning, as also implemented in [24], is very convenient when the score is computed in an ascending order from $\lvert d\rvert=1$ up to $\operatorname*{\operatorname{MAXLEN}}$. However, this is only useful in the case of looking for highly infrequent itemsets. In contrast, using the Apriori principle per se is much more useful when we start looking at itemsets of length $\operatorname*{\operatorname{MAXLEN}}$ until we reach individual categorical levels; this approach is favoured when highly frequent itemsets are of interest instead. In order to illustrate this, we present a schematic representation of how the calculation of the score would proceed for a data point $\boldsymbol{x}$ consisting of three nominal levels, say $\boldsymbol{x}=(1,1,1)^{\intercal}$. This is depicted in Figure 1, where we use a tree representation to illustrate how the search for highly frequent or infrequent itemsets is being conducted.

### v3

@@ Abstract
Outlier detection is an important data mining tool that becomes particularly challenging when dealing with nominal data. First and foremost, flagging observations as outlying requires a well-defined notion of nominal outlyingness. This paper presents a definition of nominal outlyingness and introduces a general framework for quantifying outlyingness of nominal data. The proposed framework makes use of ideas from the association rule mining literature and can be used for calculating scores that indicate how outlying a nominal observation is. Methods for determining the involved hyperparameter values are presented and the concepts of variable contributions and outlyingness depth are introduced, in an attempt to enhance interpretability of the results. The proposed framework is evaluated on both synthetic and publicly available data sets, demonstrating comparable performance to state-of-the-art frequent pattern mining algorithms and even outperforming them in certain cases. The ideas presented can serve as a tool for assessing the degree to which an observation differs from the rest of the data, under the assumption of sequences of nominal levels having been generated from a Multinomial distribution with varying event probabilities.
@@ 1 Introduction
Outlier detection (also known as anomaly detection) is a crucial aspect of data analysis that has received increasing attention by data mining researchers. The wide range of uses of outlier detection in various domains, such as in cybersecurity [1, 19], sociology [39], healthcare [45], and finance [35], has driven the development of numerous algorithms designed to detect anomalies within a data set. Most of these algorithms are restricted to just one data type and that is continuous data. However, categorical data, that is binary, ordinal and nominal data [23], is prevalent in real-world data sets, yet very few outlier detection methods can handle this data type and authors do not always make the distinction between purely nominal and ordinal variables. Outliers in data sets with nominal variables may indicate a serious offence or a critical situation. For instance, unusual combinations of symptoms, diagnoses and demographic characteristics of patients could be indications of rare conditions in healthcare monitoring or uncommon combinations of claim types, policies and accident locations may hint fraud in insurance claims. This demands the development of outlier detection methods tailored for nominal data.
@@
One approach to outlier identification assuming an underlying discrete distribution, is extending ideas from the continuous to the discrete domain. This would typically require robust estimation of the parameters of some assumed underlying distribution, computation of robust distances and finally comparing the distances to a cutoff value [see 43, 37, 38, for examples of robust estimators]. The cutoff value is typically the $1-\alpha$ quantile of the distribution of the distances, assuming validity of the hypothesised data generating mechanism. In the discrete distribution setting, this is generally a hard task that is not possible unless some model is assumed to have generated the counts in a contingency table. Log-linear models are widely used for this purpose and one can then infer what the expected proportions should be in a robust way [see for example 28, 29, 11]. However, these methods can become impractical when the number of nominal variables is very large and the analysis of a high-dimensional contingency table is required. The sparsity of the resulting contingency tables increases together with the dimensionality, which is something that needs to be accounted for as well.
@@
Another notable approach to outlier detection for nominal data is using proximity-based methods. These methods seek to detect outliers by generating a distance matrix between observations and finding the subjects that are furthest away from the rest of the points in the data set. Examples of such methods include [6] and [32]. However, distances in the categorical space often lack meaningful interpretation and implicitly make an equidistance assumption. As an example, the Hamming distance treats all mismatches equally. [42] introduced a distance metric for detecting outliers in categorical data; however, its application is restricted to $2\times 2$ contingency tables. Furthermore, these methods are very impractical in the presence of a large number of nominal variables, leading to a high computational cost.
Information-theoretic methods have also been proposed for outlier identification of nominal data [31, 47, 48]. Such methods rely on some information-theoretic measure of data disorder, such as the entropy, and typically employ a local search algorithm to optimise an objective function. As a result, these methods rely heavily on heuristics and the optimisation process can be very costly. These problems are also common in recently proposed deep learning methods [for example 13, 14].
@@
One different approach to anomaly detection for nominal data is the rule-based approach that makes use of ideas from the association rule mining literature [4]. To the best of our knowledge, the first algorithm that adopted this mindset for dealing with the problem of outlier detection in nominal data was the Link-based Outlier and Anomaly Detection in Evolving Data sets (LOADED) algorithm of [20]. In fact, LOADED was developed for identifying anomalies in mixed-type data (that is for a combination of nominal and continuous features). An improvement of this algorithm in terms of computational cost was then presented by [36], with further ameliorations being proposed by [25] for the nominal outlier identification part. More precisely, [25] introduced the Outlier Detection for Mixed Attribute Datasets (ODMAD) algorithm. The primary concern with the way ODMAD deals with nominal features is that it treats outliers in the nominal domain in a way that disregards the dimension of the subspace in which these are found. Furthermore, the implementation of ODMAD involves certain threshold parameters, the values of which need to be defined by the user, without providing guidelines of how these could be chosen in a data-driven manner.
Similarly, several frequent pattern mining methodologies have been employed for quantifying nominal outlyingness, such as the Frequent Pattern Outlier Factor (FPOF) [22] and the Frequent Pattern Isolation (FPI) [27] algorithms. However, these methods also suffer from a lack of intuition regarding the choice of their associated hyperparameter values and do not take into account that longer sequences are inherently less frequent than their subsets. Moreover, they completely overlook whether an event is rare by nature. For instance, consider a rare disease that affects only a small proportion of the population but is known and not particularly noteworthy. Frequent pattern mining algorithms will most likely flag subjects suffering from that disease as potential outliers solely due to their low frequency within a data set of clinical records, even though these are completely normal and expected cases. This underscores the need for a framework that can incorporate such contextual information and produce results that align with the underlying ground truth.
@@
The rest of the paper is organised as follows: we define the concept of nominal outlyingness by adopting a rule-based mindset in Section 2 and we give a brief introduction to key concepts from the area of association rule mining. We then propose a score that quantifies abnormal behaviour in Section 3. Our main contribution is the formulation of a score that takes into account any contextual information, relative proximity of a nominal observation to infrequency, and sparsity of the contingency tables involving multiple nominal features. Based on these considerations, we provide guidelines on the selection of the associated hyperparameters and introduce two concepts that can be used for meta-analysis purposes; these are described in Section 3 as well. Some mathematical properties of our proposed score are included in Section 4. The efficacy of our method is illustrated via a series of simulations on synthetic and publicly available data sets in Section 5, where our proposal is benchmarked against two state-of-the-art frequent pattern mining algorithms. We finally make some concluding remarks in Section 6.
@@ 2 Outlyingness in discrete space
In this Section we extend concepts of multivariate outlier detection for continuous data to discrete spaces. We define the concept of nominal outlyingness and give a brief introduction to basic notions of association rule mining.
@@ 2.1 Definition of nominal outlyingness
Outlyingness for continuous data is typically defined based on assumptions related to the data generating mechanism [18, 7], or the location of each observation relative to the rest of the points in the feature space [e.g. 10, 33]. However, location and distance are not meaningful criteria for detecting anomalies among nominal features. Assuming an underlying discrete distribution, we define outliers as values with unusually high or low frequency, analogous to low-density points in the continuous case.
Notice that looking at both highly frequent and highly infrequent values is rather unsound; for instance, a highly infrequent level of a binary variable renders the other level highly frequent by default. Therefore, one needs to specify whether their interest lies in highly frequent or highly infrequent values, that is, what they wish to define as a nominal outlier. This leads us to the following definition of nominal outlyingness:
@@ 2.2 Association rule mining background
Association rule mining is a method used to discover relationships of interest between sets of data items, with the seminal work of [4] having been applied in fields like market basket analysis [46] and customer segmentation [40]. We present some basic notions used in association rule mining.
An itemset is a set of items which occur together. Here, item refers to the value any variable takes for each observation , that is, a sequence of categorical levels in our setting. We denote an itemset by $d$ and its length by $\lvert d\rvert$. Notice that $1\leq\lvert d\rvert\leq p$, where $p$ is the number of variables and $\lvert d\rvert=1$ refers to one level of a single nominal variable. Finally, we define the support of a an itemset $d$, denoted by $\operatorname{supp}(d)$, as the number of times the itemset $d$ appears in the data.
Given a large data set, some itemsets may appear more often than others. This motivates the definition of a notion of frequency. An itemset can be considered as frequent when its support exceeds a minimum support threshold value, typically denoted by $\sigma$; otherwise it is infrequent. This threshold value may differ according to the itemset that is being looked at, therefore we use the notation $\sigma_{d}$ instead, to denote that the threshold value is specific to an itemset $d$. Determining $\sigma_{d}$ is one of the most challenging problems in the association rule mining literature; this is something that we deal with later in the paper.
Finally, we introduce the Apriori principle and the related concept of support-based pruning. According to the Apriori principle [3], if an itemset is frequent, then all of its subsets must be frequent as well. Equivalently, if an itemset is infrequent, then all its supersets are also infrequent. A direct implication of the Apriori principle is that one may not consider subsets or supersets of itemsets that are already found to be frequent or infrequent, respectively, something known as support-based pruning. The reason why support-based pruning is used is to avoid redundant computations and hence reduce the computational complexity of the search by trimming the space of itemsets that are being examined.
@@ 3 Proposed framework
The problem of outlier identification is generally complex in the sense that it is an unsupervised problem and thus, we can not be certain of whether an observation is truly outlying. However, what we can do is measure outlyingness with respect to a given scheme. By doing so, not only can we deduce how likely an observation is to be an outlier, but we can also compare observations in terms of their outlyingness. In this Section, we introduce a score that quantifies nominal outlyingness given Definition 1.
@@ 3.1 Score of nominal outlyingness
Let us assume that we are working with a dataset $\mathcal{D}$ consisting of $p$ nominal variables $\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{p}$. We denote the number of levels for each variable $\boldsymbol{X}_{i}$ by $\ell_{i}$ and assume without loss of generality that its levels are encoded by the natural numbers $1,\ldots,\ell_{i}$ $(i=1,\ldots,p)$. We treat each entry in a contingency table as an individual itemset; the total number of contingency tables is given by $2^{p}-1$ and the total number of itemsets that one can encounter is $\prod_{i=1}^{p}\left(\ell_{i}+1\right)-1$. We further define $\boldsymbol{\pi}_{i}$ to be the $\ell_{i}$-dimensional vector including the expected proportions for each level of $\boldsymbol{X}_{i}$ $(i=1,\ldots,p)$.
@@
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin(\sigma_{d},n],\\ \lvert d\rvert\leq\operatorname*{L_{\text{max}}}\end{subarray}}\frac{\sigma_{d}}{\operatorname{supp}(d)\times\lvert d\rvert^{r}},$ (1)
@@
if we are interested in highly infrequent itemsets. For highly frequent itemsets, the score of nominal outlyingness becomes
$s(\boldsymbol{x}_{i})=\sum_{\begin{subarray}{c}d\subseteq\boldsymbol{x}_{i}:\\ \operatorname{supp}(d)\notin[0,\sigma_{d}),\\ \lvert d\rvert\leq\operatorname*{L_{\text{max}}}\end{subarray}}\frac{\operatorname{supp}(d)}{\sigma_{d}\times\left(\operatorname*{L_{\text{max}}}-\lvert d\rvert+1\right)^{r}},$ (2)
@@
For the rest of this section, we assume that nominal outliers are highly infrequent observations with respect to an underlying discrete distribution. Hence, our arguments involve Expression (1); extending these to the case of highly frequent itemsets is straightforward.
The formulation in Expression (1) augments the score for observations possessing items of length up to $\operatorname*{L_{\text{max}}}$ [this is alternative notation to what 26, call MAXLEN] with support outside $(\sigma_{d},n]$ proportionally to $\sigma_{d}/\operatorname{supp}(d)$. This improves upon [26] and [25] by accounting for relative proximity to infrequency. For example, two itemsets may occur equally often, but if one is expected to be much more common, the ratio $\sigma_{d}/\operatorname{supp}(d)$ assigns it a higher degree of outlyingness.
Division by the itemset length ensures that naturally rarer, longer itemsets contribute less to the score than shorter ones, while the exponent $r$ tunes how sharply this length penalty is applied. In practice, this enables unusual single attributes (e.g. a rare occupation or diagnosis code) to provide clearer evidence of outlyingness than a rare combination of many attributes, which may simply be a consequence of combinatorial sparsity.
Despite not being explicitly mentioned in Expression (1), support-based pruning is implemented to exclude supersets of any highly infrequent itemsets from subsequent computations. That is, if $\lvert d_{1}\rvert<\operatorname*{L_{\text{max}}}$ and $\operatorname{supp}(d_{1})<\sigma_{d_{1}}$, then any itemset $d_{2}\supset d_{1}$ is not checked in subsequent calculations. We illustrate support-based pruning for a data point $\boldsymbol{x}$ consisting of three nominal levels, say $\boldsymbol{x}=(1,1,1)^{\intercal}$, in Figure 1. We use a tree representation to illustrate how the search for highly frequent or infrequent itemsets is being conducted.